OpenStudy (anonymous):

If vector A = (6.00 i -8.00 j) units, vector B = (-8.00 i + 3.00 j) units, and vector C = (25.0 i + 22.0 j) units, determine a and b such that a vector A + b vector B + vector C= 0.

OpenStudy (freckles):

\[a(6,-8)+b(-8,3)+(24,22)=(0,0) \\ 6a-8b+24=0 \text{ look at first components } \\ -8a+3b+22=0 \text{ look at second components }\] you have a system of equations to solve